OpenStudy (anonymous):

Fan and medal! Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (gorv):

3rd one both the equation are same just first one multiplied by -2

OpenStudy (gorv):

so it will produce infinite solution

OpenStudy (anonymous):

Thanks :)

OpenStudy (gorv):

4x - 3y = 9 4x=9+3y x=(9+3y)/4 put it in other equation -8((9+3y)/4) + 6 y= -18 -2(9+3y)+6y=-18 -18-6y+6y=-18 0=0 it does not depend on value of x or y so we can take any value that is infinite solution